This question is more for confirmation because I can't afford to get this wrong. The problem I have is: The number of paramecia increased exponentially. At first there were 400,000. Three days later there were 1,200,000. How many paramecia would there be at the end of 8 days? My answer is 7,720,000. Am I right? If not, would you be so kind as to teach me how to solve the problem correctly? Thank you so much!
Answer by Boreal(15235)

y=ab^x
400,000=ab^0 at day 0, and a=400,000
1200000=400000*b^3
b^3=3
b=3^(1/3)
at 8 days, y=400000*3^(1/3)^8=400000*3^(8/3)=7,488,302 rounded to nearest one.
===================
P=Poe^rt
400000=400000e^r(0);
1200000=400000e^3r
e^3r=3
ln both sides and 3r=ln 3
r=ln3/3
at 8 days
y=400000*e^8ln 3/3=7488302